# 1]organization=Scuola Normale Superiore and INFN, addressline=Piazza dei Cavalieri 7, city=Pisa, postcode=56126, state=PI, country=Italy

### v1

@@ Abstract
We explore the phenomenology of the “Dark Walker”—an $\text{SU}(3)$ theory with eight flavors of fundamental fermions in the dark sector. During inflation, its walking dynamics generate primordial non-Gaussianities through the exchange of unparticles, while accounting for the current dark matter relic abundance if the resulting massive pseudoscalars behave as strongly interacting massive particles (SIMPs). This provides a simple yet concrete example of a strongly coupled sector during inflation.
@@
In the following, we will take the Dark Walker to be an $\text{SU}(3)_{D}$ theory with $N_{f}=8$ flavors. This is an illustrator of our Dark Walker mechanism. In general, this scenario should hold for all gauge models with MFP.
@@
If we start from a weak gauge coupling in the UV, we can estimate the number of $e$-folds needed to arrive at the merged fixed point (MFP) using
@@
After inflation.—After the inflation ends, the universe becomes almost flat. The interactions $\dot{\pi}\mathcal{O}_{\psi}$ and $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\psi}$ will freeze, since the energy scale becomes too small compared to the EFT of inflation scale $M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}$. However, even though the linear mixing interaction $\dot{\pi}\mathcal{O}_{\psi}$ remains small before it gets frozen, its nature as a relevant operator will drag our model into the gapped phase, creating bound states. They are natural dark matter candidates. We expect them to be standard SIMP-like dark matter. When we take the ratio between the dark matter mass and the freeze-out temperature to be the standard choice $x_{f}\equiv m_{\text{DM}}/T_{f}\sim 20$, we can believe that the lightest state dominates. By taking the freeze-out temperature to be below the QCD phase transition scale, the dark sector will remain in thermal equilibrium with the SM particles before freeze-out. The masses of the lowest lying states of pseudoscalars (PSs) and vector mesons (Vs) can be read from [13, 21]. The mass of the lightest vector mesons is 2.2 times the mass $m_{\text{PS}}$ of the lightest pseudoscalars, which means that the vector mesons are not the light vector mesons discussed in [22]. The pseudoscalars will dominate the relic abundance, since the vector mesons freeze out first. We can infer the mass scale of the dark pseudoscalars from the observed relic abundance. Since the bound states come from a small relevant deformation-induced confinement, the interactions in the conformal phase have negligible effects on the properties of bound states.

### v2

@@ Abstract
We explore the phenomenology of the “Dark Walker”—an $\text{SU}(3)$ theory with eight flavors of massless fundamental fermions in the dark sector. During inflation, its walking dynamics generate primordial non-Gaussianities through the exchange of unparticles, while accounting for the current dark matter relic abundance if the resulting massive pseudoscalars behave as strongly interacting massive particles (SIMPs). This provides a simple yet concrete example of a strongly coupled sector during inflation.
@@
In the following, we will take the Dark Walker to be an $\text{SU}(3)_{D}$ theory with $N_{f}=8$ flavors of massless Dirac fermions. This is an illustrator of our Dark Walker mechanism. In general, this scenario should hold for all gauge models with MFP.
@@
If we start from a weak gauge coupling in the UV, we can estimate the number of $e$-folds needed to arrive near the merged fixed point (MFP) using
@@
After inflation.—After the inflation ends, the universe becomes almost flat. The interactions $\dot{\pi}\mathcal{O}_{\psi}$ and $(\partial_{\mu}\pi)^{2}\mathcal{O}_{\psi}$ will freeze, since the energy scale becomes too small compared to the EFT of inflation scale $M_{\text{pl}}|{\dot{H}_{\text{inf}}}|^{1/2}$. However, even though the linear mixing interaction $\dot{\pi}\mathcal{O}_{\psi}$ remains small before it gets frozen, its nature as a relevant operator will drag our model into the gapped phase, creating bound states. They are natural dark matter candidates. We expect them to be standard SIMP-like dark matter. When we take the ratio between the dark matter mass and the freeze-out temperature to be the standard choice $x_{f}\equiv m_{\text{DM}}/T_{f}\sim 20$, we can believe that the lightest state dominates. By taking the freeze-out temperature to be below the QCD phase transition scale, the dark sector will remain in thermal equilibrium with the SM particles before freeze-out. The masses of the lowest lying states of pseudoscalars (PSs) and vector mesons (Vs) can be read from [13, 21]. The mass of the lightest vector mesons is 2.2 times the mass $m_{\text{PS}}$ of the lightest pseudoscalars, which means that the vector mesons are not the light vector mesons discussed in [22]. The pseudoscalars will dominate the relic abundance, since the vector mesons freeze out first. We can infer the mass scale of the dark pseudoscalars from the observed relic abundance. Since the bound states come from confinement, induced by a small relevant deformation, the information of interactions in the conformal phase has negligible impacts on the properties of bound states. It is worth mentioning that if the fermions have light masses in the UV [23, 24], the DM could become the dilaton forbidden dark matter, discussed in [25].